### v3

@@ Abstract
Molecular dynamics (MD) simulations are essential for understanding atomic-scale behaviors in materials science, yet writing LAMMPS scripts remains highly specialized and time-consuming tasks. Although LLMs show promise in code generation and domain-specific question answering, their performance in MD scenarios is limited by scarce domain data, the high deployment cost of state-of-the-art LLMs, and low code executability. Building upon our prior MDAgent, we present MDAgent2, the first end-to-end framework capable of performing both knowledge Q&A and code generation within the MD domain. We construct a domain-specific data-construction pipeline that yields three high-quality datasets spanning MD knowledge, question answering, and code generation. Based on these datasets, we adopt a three stage post-training strategy—continued pre-training (CPT), supervised fine-tuning (SFT), and reinforcement learning (RL)—to train two domain-adapted models, MD-Instruct and MD-Code. Furthermore, we introduce MD-GRPO, a closed-loop RL method that leverages simulation outcomes as reward signals and recycles low-reward trajectories for continual refinement. We further build MDAgent2-RUNTIME, a deployable multi-agent system that integrates code generation, execution, evaluation, and self-correction. Together with MD-EvalBench proposed in this work, the first benchmark for LAMMPS code generation and question answering, our models and system achieve performance surpassing several strong baselines. This work systematically demonstrates the adaptability and generalization capability of large language models in industrial simulation tasks, laying a methodological foundation for automatic code generation in AI for Science and industrial-scale simulations. URL: https://github.com/FredericVAN/PKU_MDAgent2
@@ 1 Introduction
Figure 1: Overall workflow of the proposed MDAgent2, integrating data construction, model training, multi-agent runtime, and benchmark.
@@ 2.3 Evaluation of QA Ability
Table 1: Performance comparison across Overall Avg, MD-KnowledgeEval, and LAMMPS-SyntaxEval. Total Score denotes the aggregated QA performance over all question types, while scores under By Question Type correspond to different QA categories. $\Delta$ indicates the absolute improvement over the 8B baseline (Qwen3-8B). Boldface highlights the best result in each column.
Total Score By Question Type
MODEL Size / Access All $\Delta$ vs 8B Single Multi Fill OpenQA
Overall Avg
Qwen3-max Large / Closed 82.49 +11.99 91.84 72.83 77.80 90.67
Qwen3-32b 32B / Open 77.34 +6.84 89.63 69.81 65.52 88.24
@@
Qwen-flash Large / Closed 73.47 +2.97 86.87 60.84 62.17 88.45
Qwen3-14b 14B / Open 72.91 +2.41 87.46 60.93 59.43 89.04
[…]
MD-KnowledgeEval
Qwen3-max Large / Closed 86.57 +11.42 93.38 75.29 85.66 93.33
Qwen3-32b 32B / Open 81.94 +6.79 88.08 72.94 74.63 92.25
Qwen-flash Large / Closed 78.64 +3.49 89.40 61.18 71.96 93.42
Qwen3-14b 14B / Open 77.90 +2.75 87.42 63.53 68.81 92.43
[…]
LAMMPS-SyntaxEval
Qwen3-max Large / Closed 78.40 +12.56 90.30 70.37 69.94 88.01
Qwen3-32b 32B / Open 72.74 +6.90 91.18 66.67 56.41 84.23
[…]
Qwen-flash Large / Closed 68.30 +2.46 84.33 60.49 52.37 83.48
Qwen3-14b 14B / Open 67.92 +2.08 87.50 58.33 50.04 85.65
@@ 3.1 Construction of Training Datasets
MD-InstructQA: A Domain-Specific Instruction–Answer Dataset for Molecular Dynamics
@@ 3.2.2 MD-GRPO
GRPO [shao2024deepseekmath], a widely adopted algorithm for reinforcement learning in LLMs [zhang2025surveyreinforcementlearninglarge], has already demonstrated substantial value, with domain-specific variants such as Med-R1 [lai2025med] and QoQ-Med [dai2025qoq] achieving strong performance in vertical applications. Building upon GRPO, we develop a reinforcement learning framework named MD-GRPO to train LLMs to generate executable and goal-achieving LAMMPS scripts from natural-language task descriptions. The overall process consists of three stages: first, the model generates an initial code candidate based on the input task description; second, the script is executed through a scheduler to perform the molecular dynamics simulation and produce the corresponding results; finally, an evaluator model assesses the quality of the simulation outputs and returns a reward signal. This reward serves as the feedback signal for policy optimization under the GRPO framework, forming a closed-loop learning process that continuously improves both task completion and code validity.
To further enhance the model’s self-correction capability, MD-GRPO introduces a failure-feedback-tracking memory mechanism into each training iteration. Whenever the generated code fails to execute or receives a low evaluation score, the system records the underlying failure cause and reconstructs the task context accordingly. The modified context is then used as an additional training sample in the next iteration, guiding the model to avoid the previously encountered mistakes. In this way, the framework extends traditional single-turn RL into a multi-turn, trajectory-aware RL process that more faithfully simulates real runtime behavior, where multiple rounds of code refinement are often required. Consequently, the model not only learns how to produce valid simulation code, but also gradually understands why earlier attempts failed, thereby achieving an automated, reward-driven self-improvement loop.
Figure 3: Overview of the MD-GRPO training framework. The policy LLM generates candidate codes through the rollout module, which are evaluated by a reward model against a reference LLM. Group computation aggregates rewards and updates the policy via KL-regularized optimization. Low-scoring tasks are rewritten for robustness.
@@ Appendix A Tools Details
Appendix B Eval Details

### v4

@@ Abstract
Molecular dynamics (MD) simulations are essential for understanding atomic-scale behaviors in materials science, yet writing LAMMPS scripts remains highly specialized and time-consuming tasks. Although Large Language Models (LLMs) show promise in code generation and domain-specific question answering, their performance in MD scenarios is limited by scarce domain data, the high deployment cost of state-of-the-art LLMs, and low code executability. Building upon our prior MDAgent, we present MDAgent2, the first end-to-end framework capable of performing both knowledge Q&A and code generation within the MD domain. We construct a domain-specific data-construction pipeline that yields three high-quality datasets spanning MD knowledge, question answering, and code generation. Based on these datasets, we adopt a three stage post-training strategy—continued pre-training (CPT), supervised fine-tuning (SFT), and reinforcement learning (RL)—to train two domain-adapted models, MD-Instruct and MD-Code. Furthermore, we introduce MD-GRPO, a closed-loop RL method that leverages simulation outcomes as reward signals and recycles low-reward trajectories for continual refinement. We further build MDAgent2-RUNTIME, a deployable multi-agent system that integrates code generation, execution, evaluation, and self-correction. Together with MD-EvalBench proposed in this work, the first benchmark for LAMMPS code generation and question answering, our models and system achieve performance surpassing several strong baselines. This work systematically demonstrates the adaptability and generalization capability of large language models in industrial simulation tasks, laying a methodological foundation for automatic code generation in AI for Science and industrial-scale simulations. URL: https://github.com/FredericVAN/PKU_MDAgent2
keywords:
large language models, molecular dynamics, code generation, lammps
@@ 1 Introduction
Figure 1: Overall workflow of the proposed MDAgent2 framework, integrating data construction, model training, multi-agent runtime, and evaluation.
@@ 2.3 Evaluation of QA Ability
Table 1: Performance comparison across three benchmarks. Total Score denotes the aggregated QA performance over all question types. $\Delta$ indicates the absolute improvement over the 8B baseline (Qwen3-8B). Boldface highlights the best result in each column.
MODEL Size / Access Overall Avg Knowledge Syntax
All $\Delta$ All $\Delta$ All $\Delta$
Qwen3-max Large / Closed 82.49 +11.99 86.57 +11.42 78.40 +12.56
Qwen3-32b 32B / Open 77.34 +6.84 81.94 +6.79 72.74 +6.90
@@
Qwen-flash Large / Closed 73.47 +2.97 78.64 +3.49 68.30 +2.46
Qwen3-14b 14B / Open 72.91 +2.41 77.90 +2.75 67.92 +2.08
@@ 3.1 Construction of Training Datasets
MD-InstructQA: A Domain-Specific Question–Answer Dataset for Molecular Dynamics
@@ 3.2.2 MD-GRPO
GRPO [shao2024deepseekmath], a widely adopted algorithm for reinforcement learning in LLMs [zhang2025surveyreinforcementlearninglarge], has already demonstrated substantial value, with domain-specific variants such as Med-R1 [lai2025med] and QoQ-Med [dai2025qoq] achieving strong performance in vertical applications. Building upon GRPO, we develop a reinforcement learning framework named MD-GRPO to train LLMs for generating executable and goal-achieving lammps scripts from natural-language task descriptions. Given an input query, the policy model produces multiple candidate codes, which are executed in a molecular dynamics simulation environment and evaluated by a reward model based on code and results.Then policy optimization, forming an execution-driven closed-loop learning process.
To further enhance the model’s self-correction capability, MD-GRPO introduces a failure-aware trace-and-rewrite mechanism(Trace Module in Fig. 3) into each training iteration. When a generated code fails to execute or receives a low evaluation score, the system records the entire trajectory, including the low-scoring code and its corresponding evaluation feedback, and rewrites the original q into a refined form q’. The rewritten query is then added to the training data pool, enabling the model to iteratively correct previously observed failure patterns.
Through this design, the framework extends conventional single-step reinforcement learning to trajectory-aware optimization that more closely reflects real runtime behavior, where multiple rounds of code refinement are often necessary. As a result, the model not only learns how to generate valid simulation code, but also progressively understands why earlier attempts failed and how to correct them, forming an automated, reward-driven self-improvement loop.
Figure 3: Overview of the MD-GRPO. The framework integrates code generation, simulation execution, and trajectory rewriting to enable execution-driven optimization for molecular dynamics tasks.
@@ 4 Discussion
Finally, while this work focuses on molecular dynamics as a representative domain, the proposed data construction paradigm, execution-aware reinforcement learning, and agentic closed-loop workflow are broadly applicable to other AI-for-Science scenarios involving simulation-driven reasoning and code generation.
@@ Appendix A Tools Details
Appendix B Evaluation Details